OpenStudy (anonymous):

A square poster has sides measuring 2 feet less than the signs of a square sign. If the difference between the two areas is 12 square feet, find the lengths of the sides of the poster and the sign.

OpenStudy (anonymous):

If the square sign has a side of "x" then The square poster has sides of "x-2" x^2 - (x-2)^2= 12 x^2 - (x^2 - 4x + 4) = 12 x^2 - x^2 + 4x - 4 = 12 4x - 4 = 12 4x = 16 x = 4 Square sign is 4 x 4 Square poster is 2 x 2 16 - 4 = 12